OpenStudy (anonymous):

The mass of the Earth is 5.9742 x \[10^{24}\] The mass of the sun is 333 000 times the mass of the Earth Work out the mass of the Sun Give your answer in the standard form correct to 3 significant figures Can someone help me on this please?

OpenStudy (anonymous):

Well, if the mass of the earth is \(5.9742\times10^{24}\) and the mass of the sun is 333,000 times that.. What is 333,000 times \(5.9742\times10{^24}\)

OpenStudy (anonymous):

err 333,000 times \(5.9742\times10^{24}\)

OpenStudy (anonymous):

Cause that's the mass of the sun.

OpenStudy (anonymous):

i typed it into my calculator and it came up with 1.9894419 x \[10^{30}\]

OpenStudy (anonymous):

Seems reasonable. 5.97*3.33 = 19.88, and \(10^5 \times 10^{24} =10^29\). Moving the decimal one more place would be \(1.988 \times 10^{30}\)

OpenStudy (anonymous):

And I rounded down, so that's probably correct.

OpenStudy (anonymous):

tyvm